OpenStudy (anonymous):

help me please if t=tanh(1/2 x), prove sinhx=2t/(1-t^2)

OpenStudy (anonymous):

thank you, so here's the question, isnt tanhx=\[\frac{ e^x - e^{-x} }{ e^x + e ^{-x} }\]

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle t=\frac{e^{(x/2)}-e^{-(x/2)}}{e^{(x/2)}+e^{-(x/2)}} }\) \(\color{#000000 }{ \displaystyle t=\frac{e^{x}-1}{e^{x}+1} }\) We can solve for \(e^x\) in terms of \(t\). \(\color{#000000 }{ \displaystyle t=\frac{e^{x}-1}{e^{x}+1} }\) \(\color{#000000 }{ \displaystyle e^x=\frac{1+t}{1-t} }\) \(\color{#000000 }{ \displaystyle t=\frac{e^{x}-1}{e^{x}+1} \quad \Longrightarrow \quad \frac{1+t}{1-t}=e^{x}\quad \Longleftrightarrow \quad \frac{1-t}{1+t}=e^{-x} }\) then, \(\color{#000000 }{ \displaystyle \sinh(x)=\frac{e^{x}-e^{-x}}{2} =\frac{2t}{1-t^2} }\) \(\color{#000000 }{ \displaystyle \frac{\dfrac{1+t}{1-t}-\dfrac{1-t}{1+t}}{2} =\frac{2t}{1-t^2} }\)

OpenStudy (solomonzelman):

Sorry for making that erro, but nice that you caught it.

OpenStudy (solomonzelman):

and yes, it becomes indeed equivalent when algerba correctly applied.

OpenStudy (anonymous):

Ok, tq, I'll try first